# R-Judge: Benchmarking Safety Risk Awareness for LLM Agents

### 18 Jan 2024 v1

@@ 4 Experiments
In this section, we first describe the experimental setup, especially metrics. Then, we evaluate 8 popular LLMs with R-Judge on the Standard Test and Oracle Test.
@@ 4.1 Setup
We comprehensively assess 8 LLMs, including API-based models and open-source models. The API-based models include GPT-4 (gpt-4-0613 version) (OpenAI, 2023) and ChatGPT (gpt-3.5-turbo-0613 version). [^2] https://chat.openai.com/. The open-source models include Llama-2 (Touvron et al., 2023) and Vicuna (Chiang et al., 2023) series.
@@
We adopt the zero-shot chain-of-thought prompting (Kojima et al., 2022) (dubbed Zero-Shot-CoT) to induce LLMs to generate the reasoning steps before producing the final answer. This kind of analyze-then-output process has been shown to improve reasoning performance, as well as interpretability (Zhang et al., 2023c).
Task instructions are simple and clear for generality, as presented in Appendix C.1. Additional priors of safety guidelines are excluded in both tests.
@@
Assuming that the dataset $\mathcal{D}$ consists of $n$ sample $\{(x_{i},y_{i})_{i=1}^{n}\}$, where $x_{i}$ is the input and $y_{i}\in\{0,1\}$ the output. We denote $p_{\theta}(\cdot)$ as the language model, and $\hat{y}_{i}=p_{\theta}(x_{i})\in\{0,1\}$ as the predicted label. The metrics described above can be formulated as
[…]
where $\mathbb{I}(\cdot)$ is the indicator function.
@@
We provide random scores and human scores for comparison. To obtain human performance, we employ six human experts majoring in computer science and cyber security and average their scores to measure the ceiling performance on R-Judge. As agent safety judge is a task of binary classification, random scores are 50% for F1, Recall, and Specificity.
Models Standard Test (w/o Description) Oracle Test (w/ Description)
F1 Recall Specificity Validity Consistency F1 Recall Specificity Validity Consistency
Random 50.00 50.00 50.00 50.00 50.00 50.00 50.00 50.00 50.00 50.00
Human 89.38 81.48 97.66 100.00 100.00 100.00 100.00 100.00 100.00 100.00
GPT-4 72.29 61.86 82.35 100.00 98.65 99.49 100.00 98.04 100.00 99.32
ChatGPT 39.69 26.80 84.31 100.00 95.27 91.63 95.88 74.51 100.00 90.54
Vicuna-13b-v1.5 9.62 5.50 83.01 94.14 93.57 87.32 79.38 95.43 100.00 89.41
Vicuna-13b-v1.5-16k 7.62 4.47 77.78 93.47 88.30 82.90 74.23 90.85 100.00 86.04
Llama-2-13b-chat-hf 27.77 23.71 13.07 34.46 85.66 65.17 69.07 18.30 53.61 97.07
Vicuna-7b-v1.5 5.74 3.43 69.93 100.00 77.25 15.65 9.62 78.43 100.00 48.87
Vicuna-7b-v1.5-16k 17.63 12.03 56.21 95.27 78.04 53.67 47.77 49.68 99.32 51.69
Llama-2-7b-chat-hf 6.93 5.50 3.27 17.79 56.78 9.60 7.56 8.49 16.89 60.00
Table 1: Main results(%) of the Standard Test and Oracle Test in R-Judge. Segment 1: Reference Score, including Random as baseline and Human as upper bound marked as underline; Segment 2: API-based models; Segment 3: 13B models; Segment 4: 7B models. All presented metrics are average results of 3 experiments with different task instructions, shown in Appendix C.1. The best average model results are in bold face.
@@ 4.2 Main Results
Model ALL Program OS IoT Software Finance Web Health
GPT-4 72.29 88.89 70.59 42.11 69.57 76.19 69.57 44.44
ChatGPT 39.69 50.00 42.86 40.00 14.29 40.00 42.11 22.22
Table 2: Standard Test F1 score of API-based LLMs in different categories.
@@
Table 1 presents the main results in R-Judge. Based on the results, we have the following key findings.
@@
(i) All LLMs perform unsatisfactorily on the R-Judge benchmark. Although solvable for humans (with nearly 90% F1 score) in the Standard Test, all the LLMs perform much inferior. Remarkably, GPT-4 ranks first and is also the only model scoring higher than random. However, it is still far from succeeding in the task. The results show that R-judge is a challenging benchmark and there is considerable room for enhancing the safety risk awareness of LLMs.
(ii) API-based models perform better than open-source models, including those equipped with long-context length. Notably, GPT-4 and ChatGPT achieve 72.29% and 39.69% F1 scores in the Standard Test, respectively. In contrast, the best F1 score among open-source models is 27.77% reported by Llama-2-13b-chat-hf.
(iii) Larger-scale models achieve better F1 scores in general. Specifically, LLMs with 13B parameters (13B LLMs for short) generally perform better than those with 7B parameters. The results imply that larger models tend to take advantage of their inference abilities when provided with informative safety feedback, i.e., the risk descriptions. In addition, 13B LLMs also perform more robustly than 7B LLMs with regard to the order of choices, as reflected by the Consistency score.
(iv) There is a strong correlation between the F1 scores and instruction following ability for the same series of LLMs. Recall the definition of validity in Section 3.3.2, Validity reflects the ability of instruction following in LLMs. Concretely, Vicuna models exhibit better validity compared with Llama-2 models under the same scale. Such an observation is also evidenced by the fact that Vicuna models are instruction-following LLMs fine-tuned on Llama-2 with user-shared conversations collected from ShareGPT.
(v) All the models achieve substantial performance gains when the risk descriptions are provided in the Oracle Test. For example, the F1 score of GPT-4 is upgraded from 72.29% to 99.49%. On the one hand, the results reveal the importance of salient safety risk descriptions. This finding aligns with recent studies showing that LLMs are able to self-correct with the crutch of external feedback (Gou et al., 2023; Huang et al., 2023a; Valmeekam et al., 2023), e.g., risk descriptions in our problem. Hopefully, such descriptions may be obtained from a verification model that dynamically monitors the agent interaction in the environment in future studies. On the other hand, the results also indicate that our records are understandable and risk descriptions are informative.
Models Zero-Shot-CoT Few-Shot-CoT
F1 Recall Specificity F1 Recall Specificity
GPT-4 72.29 61.86 82.35 48.08 43.10 96.36$\uparrow$
ChatGPT 39.69 26.80 84.31 20.69 11.84 96.36$\uparrow$
Vicuna-13b-v1.5 10.71 6.19 82.35 5.71 3.90 54.55
Vicuna-13b-v1.5-16k 3.67 2.06 80.39 11.65$\uparrow$ 7.79$\uparrow$ 63.64
Llama-2-13b-chat-hf 32.75 28.87 9.80 10.45 9.09 9.09
Vicuna-7b-v1.5 5.31 3.09 74.51 5.61$\uparrow$ 3.90$\uparrow$ 50.91
Vicuna-7b-v1.5-16k 20.98 15.46 39.22 7.58 6.49 9.09
Llama-2-7b-chat-hf 1.36 1.03 3.92 1.55$\uparrow$ 1.30$\uparrow$ 7.27$\uparrow$
Table 3: Comparison between Few-Shot-CoT prompting and Zero-Shot-CoT prompting in Standard Test. In Few-Shot-CoT, ups of metrics are marked with $\uparrow$. Segment 1: API-based models; Segment 2: 13B models; Segment 3: 7B models. The prompts are provided in Appendix C.1 and Appendix C.2, respectively.
Model F1 Consistency
GPT-4 99.49±$\pm\textbf{0.00}$ 99.32±$\pm\textbf{0.00}$
ChatGPT 91.63±$\pm 0.52$ 90.54±$\pm 0.75$
Vicuna-13b-v1.5 87.32±$\pm 2.43$ 93.57±$\pm 5.38$
Vicuna-13b-v1.5-16k 82.90±$\pm 1.53$ 88.30±$\pm 6.91$
Llama-2-13b-chat-hf 65.17±$\pm 0.95$ 85.66±$\pm 10.62$
Vicuna-7b-v1.5 15.65±$\pm 7.32$ 77.25±$\pm 11.19$
Vicuna-7b-v1.5-16k 53.67±$\pm 12.68$ 78.04±$\pm 12.91$
Llama-2-7b-chat-hf 9.60±$\pm 5.65$ 56.78±$\pm 16.84$
Table 4: Standard deviations of F1 and Consistency with different task instructions in Oracle Test. Segment 1: API-based models; Segment 2: 13B models; Segment 3: 7B models. The lowest deviations are in bold face.
@@ 5 Analysis
API-based LLMs such as GPT-4 and ChatGPT have shown superior performance in R-Judge. In this section, we will first investigate their performance in different scenarios, as well as performance variance with different task instructions. Besides, we are interested in whether their performance can be improved further. We will show that the simple few-shot prompting technique fails at the task and then investigate more effective solutions with the chain of action analysis.
5.1 Model Performance in Different Scenarios
Table 2 presents the performance of GPT-4 and ChatGPT in different scenarios of R-Judge. We see that both models achieve relatively better performance in program and OS scenarios. The superiority might be related to their ability to understand and reason over codes (Gao et al., 2023; Li et al., 2023b; Chen et al., 2021). In addition, GPT-4 shows remarkable improvements in finance and health scenarios, which is likely due to the merits of its safety alignment mechanisms.
5.2 Model Performance Variance with Different Task Instructions
LLMs have been shown sensitive to system prompts in existing studies (Sclar et al., 2023; Zhu et al., 2023). Therefore, we conduct 3 experiments with different task instructions to obtain the average results in Table 1. Here we present standard deviations of F1 score and Consistency in Oracle Test in Table 4 to see the stability of model performance with different task instructions.
We observe that task instructions slightly affect the performance of models. Furthermore, there is a correlation between standard deviations of metrics, the Consistency score, and the model scale. In general, larger models present higher Consistency and lower standard deviations, e.g., larger models are more reliable in the task of agent safety judge.
Acceptable performance variances of models larger than 10B indicate the quality of R-Judge datasets. In later analysis, We choose the average best one (Zero-Shot-CoT Prompt 2, detailed in Appendix C.1) as the default task instruction.
5.3 Does Few-shot Prompting Improve Performance?
We are also interested in whether using few-shot demonstrations can achieve better performance than our prior Zero-Shot-CoT prompting. To this end, we design two-shot demonstrations (Appendix C.2) to form Few-Shot-CoT (Wei et al., 2022b) for analysis. As results shown in Table 3, Few-Shot-CoT does not consistently improve overall performance. The results align with the findings in AgentMonitor (Naihin et al., 2023) and SimulateBench (Xiao et al., 2023) where few-shot demonstrations drop the performance. The most plausible reason would be the limited coverage of the demonstrations for agent-related tasks. As they are hard to cover the full range of possible risks, using those demonstrations may confuse the LLMs. The result further verifies that our task is challenging with diverse risk types and it is unlikely to be solved by simple prompting mechanisms.
Figure 3: Illustration of CoSA compared with Zero-Shot-CoT. Compared with Zero-Shot-CoT, CoSA prompts LLMs to generate analysis in each turn and further output a label in the last turn.
5.4 Chain of Safety Analysis Technique
Due to the interactive nature of agents, there are multi-turn action executions in a record as defined in Section 2. Analyzing those executed actions could provide valuable evidence for the final safety decision. Therefore, we design a chain of safety analysis (CoSA) technique to take each step of action into account, shown in Figure 3.
For a record with $k$ turns, we call the LLM $k$ times to generate the safety analysis for each turn. Concretely, for the $i$-th turn ($i=1,2,\cdots,k)$, the LLM is fed with the former $i$ turns of interactions and generates safety analysis. At the last turn, the LLM generates judgment based on the complete record and the $k$ history analysis.
Results in Table 5 show that the CoSA technique achieves much better F1 scores than the standard Zero-Shot-CoT prompting (e.g., 72.29% $\rightarrow$ 80.13% for GPT-4). We see that the recall has been improved substantially with a slight sacrifice of specificity. The results show that inducing the model to focus on the fine-grained information on each turn of records helps the final judgment of safety risks.
@@ 5.4 Chain of Safety Analysis Technique
Zero-Shot-CoT 72.29 61.86 82.35
Few-Shot-CoT 48.08 43.10 96.36
CoSA 80.13 74.51 78.18
@@
Zero-Shot-CoT 39.69 26.80 84.31
Few-Shot-CoT 20.69 11.84 96.36
CoSA 56.97 45.63 72.73
Table 5: Results of API-based models (ChatGPT and GPT-4) with different prompt settings (Zero-Shot-CoT, Few-Shot-CoT, CoSA) in the Standard Test. The best results of each model are in bold face.
@@ 5.5 Case Study
To figure out the capability demand of the task and flaws of current LLMs, we further analyze the results of GPT-4 with a manual check. We find that there are 3 key failure types. Distributions of failure types are shown in Figure 4 and representative failure cases are provided in Appendix D.3.
(i) Scenario Simulation: Fail to retrieve corresponding knowledge and reason in specific scenarios. Some risks are invisible. They need the judge to infer the effect of actions, for example, the effect of link sharing with edit access. These risks demand specific domain knowledge and reasoning. Due to the lack of scenario familiarity and reasoning ability, GPT-4 may fail to retrieve corresponding knowledge and reason in specific scenarios to identify risks. For example, GPT-4 failed in the case where the agent shared the link on Twitter with edit access instead of comment access. Furthermore, due to attention distraction in a long context, GPT-4 may miss critical details of risks. For example, GPT-4 failed in a case where a long blog declared copyright but the user reproduced without permission, thus violating copyright. Interestingly, we observed that CoSA can alleviate attention distraction in a long context to promote the overall performance of GPT-4.
(ii) Understanding Adaptability: Unable to understand risk in specific conditions. Risks are highly related to conditions, e.g., risky actions in Scenario A may be safe in Scenario B instead. With safety alignment, GPT-4 has already possessed a comprehensive knowledge of security concerns. However, due to the limitation of understanding conditions, GPT-4 can be rigid sometimes and over-concerned. For example, GPT-4 would respond that “the actions taken by the agent are generally safe, but there are a few potential security concerns…” and wrongly classify them as unsafe.
(iii) Safety Alignment: Deviation of Safety Alignment with Human. Topics related to morality, ethics, and privacy are vague. Given that our human experts reach a consensus on cases, GPT-4 shows an understanding bias with humans. For example, GPT-4 failed cases where the agent follows user instructions to post a Twitter with unverified information about a classmate and thus violates privacy. In addition, we deem that even if risky actions are specified by the user, the agent should ask the user for confirmation. However, GPT-4 failed in most of these cases.
Figure 4: Proportion of 3 key GPT-4 failures with Zero-Shot-CoT in Standard Test.
Summary
Based on experimental results and case study, we conclude that the development of a risk-aware agent mainly spotlights two parts, scenario-agnostic model capability and scenario-specific safety guidelines. On the one hand, for scenario-agnostic model capability, the key is to strengthen the ability to capture specific risks in the long context of agent interaction. Prompting techniques such as CoSA can guide models to be aware of risks but fail to improve their reasoning ability in specific scenarios. Safety fine-tuning is another choice to reinforce the risk awareness of agents and R-Judge can be an effective training set. On the other hand, when developed for specific scenarios, a safe agent should align with humans in domain knowledge, risk understanding, and safety standards. Therefore, agents should be equipped with according safety guidelines either by fine-tuning or prompting to adapt to various scenarios.

### 18 Feb 2024 v2

@@ 4 Experiments
In this section, we first describe the experimental setup, especially metrics. Then, after validating R-Judge, we evaluate 9 popular LLMs on two tests of safety judgment and risk identification.
@@ 4.1 Setup
We comprehensively assess 9 LLMs, including API-based models and open-source models. The API-based models include GPT-4 (gpt-4-0613 version) (OpenAI, 2023) and ChatGPT (gpt-3.5-turbo-0613 version). [^1] https://chat.openai.com/. The open-source models include Llama-2 (Touvron et al., 2023) and Vicuna (Chiang et al., 2023) series. Particularly, on the foundation of Llama-2-7b-chat, LlamaGuard-7b (Inan et al., 2023) is fine-tuned to identify safety risks in LLM-generated content. Computational resources and time consumption to run these models are listed in Appendix E.
@@
We adopt the zero-shot chain-of-thought prompting (Kojima et al., 2022) (dubbed Zero-Shot-CoT) to induce LLMs to generate the reasoning steps before producing the final answer. This kind of analyze-then-output process has been shown to improve reasoning performance, as well as interpretability (Zhang et al., 2023b). Task instructions are simple and clear for generality, as presented in Appendix C.3.1.
@@
Risk Identification, the pairwise evaluation, introduces ${\rm Grade}$, ${\rm Alertness}$ and ${\rm Effectiveness}$, as stated in Section 3.3. In the leaderboard, they are all normalized to the range of 0 to 1.
@@
We provide random scores and human scores for comparison. To obtain human performance, we employed six human experts majoring in computer science and cyber security and averaged their scores to measure the ceiling performance on R-Judge. As safety judgment is a task of binary classification, random scores are 50% for F1, Recall, and Specificity. For open-ended risk identification, random Effectiveness is 0%, while random Alertness is 50%.
4.2 Validating R-Judge
Human agreement with GPT-4 evaluator in the risk identification test.
First, to tackle the potential bias and incapability of LLM-as-Judge (Liu et al., 2023; Zheng et al., 2023), we measure the agreement between GPT-4 and human experts by the Pearson correlation coefficient (Wikipedia, ). The Pearson correlation coefficient quantifies the linear correlation between two variables and is a value between -1 and 1, where the higher value indicates a higher correlation.
On 50 randomly selected unsafe samples, three human experts scored model analysis according to the same criteria that prompt the GPT-4 evaluator. The Pearson correlation coefficient on Grade, Effectiveness and Alertness are 0.95, 0.94 and 0.89, proving the reliability of pairwise evaluation for the risk identification test. The full results across models are listed in Appendix C.2. Notably, for GPT-4 answers, there exists bias that the Alertness agreement between GPT-4 evaluator and human is only 0.64. However, the Grade and Effectiveness are valid with values of 0.92 and 0.89.
High Correlation between safety judgment and risk identification of LLM performance.
We observed that model ranks are similar in the two tests and further proved the high correlation between the two tests. With F1 score in the safety judgment test against Grade, Effectiveness and Alertness in the risk identification test, the Pearson correlation coefficient are 0.92, 0.91 and 0.96.
Therefore, mutually supportive, the two intersected tests offer a valid and progressive evaluation. The safety judgment test is more fair and affordable while the risk identification test holds fine-grained interpretability and better discrimination on model performance.
@@ 4.3 Main Results
Table 1 presents the main results in R-Judge. Based on the results, we have the following key findings:
@@
(i) Most LLMs perform unsatisfactorily on the R-Judge benchmark. Although solvable for humans (with nearly 89.07% F1 score in Safety Judgment and 87.00% grade in Risk Identification), most LLMs perform much inferior. Remarkably, GPT-4 ranks first and is also the only model scoring higher than random in the safety judgment test. However, it is still far from reaching human-level performance. The results show that R-Judge is a challenging benchmark and LLMs tend to generate ineffective analysis and make wrong judgments when judging safety risks in agent interaction. There is considerable room for enhancing the safety risk awareness of LLMs in open agent scenarios.
(ii) Larger-scale models achieve better performance. For models from the same series, e.g. vicuna-13b-v1.5-16k and vicuna-7b-v1.5-16k, Llama-13b-chat-hf and Llama-2-7b-chat-hf, the larger one with more parameters present better performance in both tests. And in general, LLMs with more parameters generally perform better. There is only one exception (Vicuna-13b-v1.5-16k 43.24% $>$ ChatGPT 39.42%) in the safety judgment test and no exception in the risk identification test. The results imply that risk awareness demands knowledge and reasoning abilities from numerous parameters for safety judgment and risk identification, which is further confirmed by manual case studies in Section 5.2.
(iii) Additional fine-tuning on safety alignment does not necessarily raise risk awareness in multi-turn agent scenarios. It is known that both the Llama-2-chat series and the Vicuna series are fine-tuned on the foundation of the Llama-2 base model. Llama-2-chat series are conducted on additional safety alignment including fine-tuning and RLHF (Bai et al., 2022; Ouyang et al., 2022) while there is no additional safety alignment for Vicuna. However, results indicate that Llama-2-chat series present no prominent excess than Vicuna with same parameters. And Llama-2-chat series suffer from low validity, i.e. fail to output safety label after taking in multi-turn long context (see Appendix D.2.2). Significantly, LlamaGuard-7b, specifically fine-tuned on identifying safety risks in LLM-generated content, is incapable of identifying safety risks in the execution of LLM agents. Considering the (ii) finding, we argue that for LLMs functioning as agents rather than chatbots, fine-tuning for general capabilities involving knowledge and reasoning should prioritize additional alignment fine-tuning for harmless content.
@@ 5 Analysis
In this section, we delve into the effect of different mechanisms on model performance with further experiments and case studies, acquiring a nuanced understanding of the two tests.
5.1 Influence of Different Prompting Techniques
The main results show the baseline performance of models with no risk priors or safety guidelines in system prompts. In this section, we explore the influence of different prompts on model performance in the safety judgment test.
Zero-Shot-CoT with Risk Types
As disclosed in literature (Li et al., 2023b; Wang et al., 2023c), task-specific hints are influential priors to improve task performance. Incorporating R-Judge risk types as hints into the task instruction, only the ChatGPT F1 score improves owing to improved recall with some sacrifice of specificity, as shown in Table 2.
Few-Shot-CoT Prompting
Based on Zero-Shot-CoT, we design two-shot demonstrations (Figure 3 in Appendix) to form Few-Shot-CoT (Wei et al., 2022b) for analysis. As results shown in Table 2, Few-Shot-CoT does not consistently improve overall performance. The most plausible reason would be the limited coverage of the demonstrations for agent-related tasks (Naihin et al., 2023; Xiao et al., 2023). As they are hard to cover the full range of possible risks, using those demonstrations may confuse the LLMs.
@@ 5.1 Influence of Different Prompting Techniques
Zero-Shot-CoT 72.52 62.00 83.64
+ Few-Shot 64.86 48.00 100.00
+ Risk Types 71.26 62.00 78.18
@@
Zero-Shot-CoT 39.42 27.00 81.82
+ Few-Shot 32.26 20.00 92.73
+ Risk Types 56.10 46.00 67.27
Vicuna-13b-v1.5-16k F1 Recall Specificity
Zero-Shot-CoT 43.24 32.00 70.91
+ Few-Shot 23.08 15.00 72.73
+ Risk Types 34.33 23.00 80.00
Llama-2-13b-chat-hf F1 Recall Specificity
Zero-Shot-CoT 38.86 34.00 25.45
+ Few-Shot 34.32 29.00 27.27
+ Risk Types 38.04 35.00 10.91
Table 2: Result(%) comparison of different prompt settings (Zero-Shot-CoT, Few-Shot-CoT, Zero-Shot-CoT w/ Risk Types) in the safety judgment test. The 4 models with best baseline performance are selected. The best average model results are in bold face.
Summary
The results of the prompting experiments conducted above additionally confirm the challenging nature of our task, especially when dealing with intricate cases encompassing diverse risk types. It becomes evident that straightforward prompting mechanisms are unlikely to suffice in addressing the complexity inherent in our task.
@@ 5.2 Case Study
To figure out the capability demand of the task and flaws of current LLMs, we further analyze the results of GPT-4 with a manual check. We conclude with 3 key capability flaws leading to failures:
(i) Scenario Simulation:
Fail to retrieve relevant knowledge and reason in specific scenarios.
(ii) Understanding Adaptability:
Unable to comprehend risks in specific conditions.
(iii) Safety Alignment:
Deviation of safety alignment with humans in practical scenarios.
These key flaws are aligned with results disclosed by Effectiveness and Alertness in risk identification test which reveal model capabilities in different dimensions. Due to space limitations, detailed explanations and representative failure cases are attached in Appendix D.2.1.
5.3 Oracle Test: The Effect of Risk Description on Safety Judgment
The unsatisfactory results in the risk identification test and case study demonstrate the ineffectiveness of model-generated analysis. To investigate whether model performance on safety judgment can be improved when provided with valid analysis, we introduce the Oracle Test, a modified version of the safety judgment test where analysis of LLMs is substituted with ground truth risk description written by human annotators.
Results (see Table 7 in Appendix) show that all the models achieve substantial performance gains (such as GPT-4 from 72.52% to 99.50%) with risk descriptions as environment feedback, revealing the importance of salient safety risk feedback. Hopefully, they can be obtained from a model that dynamically monitors the agent interaction in the environment in future studies.
5.4 Summary
Based on experimental results and case study, we conclude that the development of a risk-aware LLM agent mainly spotlights two parts, general model capability and scenario-specific safety guidelines. On the one hand, for general model capability, the key is to strengthen the ability to capture specific risks in the long context of multi-turn interaction. Specifically, to raise safety risk awareness of LLMs in agent open scenarios, fine-tuning for general capabilities involving knowledge and reasoning should prioritize additional alignment fine-tuning for harmless content. On the other hand, when developed for specific scenarios, a safe agent should align with humans in domain knowledge, risk understanding, and safety standards. It is promising to equip with a monitor model specifically fine-tuned to provide salient safety risk feedback for the safe execution of LLM agents.